OpenStudy (vheah):

Please check my equations!!! If the book is initially at rest, what is its speed after it has traveled 0.400 m up the wall?

OpenStudy (vheah):

Here's the picture of the problem:

OpenStudy (vheah):

I drew a free body diagram which is this:

OpenStudy (vheah):

Here are the equations i got out of that FBD : \[\sum_{}^{}Fx = N - Fcos(60) = 0\] \[\sum_{}^{} Fy = f _{k} + F \sin (60) = ma \] I know that \[f _{k} = \mu(Fcos(60))\] can someone check my equations? Should i find acceleration and then plug it in the kinematics equation \[v^2_{f} = v^2_{i} + 2*a*d\] where inital velocity is 0?

OpenStudy (vheah):

@pooja195 can you check what I have?

OpenStudy (jiteshmeghwal9):

The only mistake u did is that u didn't count gravity while forming equation of net force along y-axis

OpenStudy (jiteshmeghwal9):

Net force along y-axis is \[F sin 60 - f_k-mg=ma\]

OpenStudy (vheah):

oh okay! Thank you! @jiteshmeghwal9